OpenStudy (anonymous):

I need to find the zeros of the function f x)=(x+5)^2 +3. Step by Step please!!!

OpenStudy (anonymous):

this has no zeros, because no matter what x is \[(x+5)^2\geq 0\] and so \[(x+5)^2+3\geq 3\]

OpenStudy (anonymous):

The answer was -5 +/- i squareroot 3. There was an imaginary number. Does that mean it uses the -b +/- squareroot b^2-4(a)(c) over 2(b)?

OpenStudy (anonymous):

when you use the quadratic equation (the one you described above) you get \[-5 \pm i \sqrt{3}\]. Those are your zeroes, although you get an imaginary number, so it is a complex number

OpenStudy (anonymous):

can you show me step by step?

OpenStudy (anonymous):

I tried it myself but it didn't work out so well.